# MM-Interleaved: Interleaved Image-Text Generative Modeling via Multi-modal Feature Synchronizer

### v1

@@ Abstract
* Equal contribution.† This work is done when Changyao Tian, Weiyun Wang, and Zhe Chen are interns at Shanghai AI Laboratory.🖂 Corresponding author (daijifeng@tsinghua.edu.cn).
Developing generative models for interleaved image-text data has both research and practical value. It requires models to understand the interleaved sequences and subsequently generate images and text. However, existing attempts are limited by the issue that the fixed number of visual tokens cannot efficiently capture image details, which is particularly problematic in the multi-image scenarios. To address this, this paper presents MM-Interleaved, an end-to-end generative model for interleaved image-text data. It introduces a multi-scale and multi-image feature synchronizer module, allowing direct access to fine-grained image features in the previous context during the generation process. MM-Interleaved is end-to-end pre-trained on both paired and interleaved image-text corpora. It is further enhanced through a supervised fine-tuning phase, wherein the model improves its ability to follow complex multi-modal instructions. Experiments demonstrate the versatility of MM-Interleaved in recognizing visual details following multi-modal instructions and generating consistent images following both textual and visual conditions. Code and models are available at https://github.com/OpenGVLab/MM-Interleaved.
@@ 1 Introduction
Interleaved image-text data, where the text and images are related in content and intertwined in layout, is very common on the internet. As an extension of image-text pairs widely used in previous multi-modal models [37, 21, 82, 22, 70], the interleaved format not only covers a broader range of data but also presents more complex information structures. Developing multi-modal models that can simultaneously understand and generate such interleaved image-text data has significant research value and practical application potential. It broadens the application scope of previous multi-modal models [83] and enables more unified processing of multi-modal data, bringing previously isolated research fields such as text-to-image generation [74, 71] and visual question-answering [48, 53] together.
Figure 1: Illustration of multi-modal feature synchronizer. In the auto-regressive generation of interleaved image-text sequences, besides interacting with low-resolution image tokens with self-attention, tokens in MM-Interleaved could also use the multi-modal feature synchronizer to cross-attend multi-scale high-resolution image features. The multi-modal feature synchronizer ensures the cross attention will have exactly the same causal relation between image and text tokens.
@@
With recent progress in multi-modal modeling with Large Language Models (LLMs) [7, 64], exploring how LLMs can be leveraged for interleaved image-text modeling has become increasingly attractive. A core challenge of interleaved multi-modal LLMs is how to effectively handle multiple images within the limited context length of LLMs. In most multi-modal LLMs, images are encoded as visual tokens from an image tokenizer (also known as image encoder) and fed into LLMs together with text tokens [43, 18, 83, 40]. To reduce computational demands and the required context length of LLMs, Perceiver Resamplers [3] are commonly used to map each image from up to 1024 tokens to a fixed number of visual tokens (e.g., 32 or 64 tokens) as in Fig. 2. Due to the relatively small token number, critical image details may be overlooked, especially in tasks requiring fine-grained observation. Increasing the number of visual tokens [55, 99] can address this issue to some extent. While the context length of LLMs is typically limited, allocating a large number (e.g., 2890 in [84]) of visual tokens per image would pose a significant limitation in multi-image scenarios.
@@
We note that the main problem is that the Perceiver Resampler only takes image features as input and is context-insensitive. It compresses each input image to a fixed number of tokens, which makes it harder or even impossible to preserve all information required and fulfill all needs for subsequent generation. In contrast, allowing intermediate layers in decoder networks to dynamically extract a small amount of relevant information from the images can compensate for this information loss. Such a dynamic extraction mechanism should also efficiently handle scenarios requiring multiple images and high-resolution feature maps.
Motivated by this, we designed a fine-grained multi-modal feature synchronizer based on the deformable sparse attention [110]. As shown in Fig. 1, it enables the multi-modal LLMs to observe multi-scale high-resolution image features from multiple images efficiently. This mechanism facilitates the extraction of fine-grained image details with only a small number of input visual tokens, which is particularly applicable to interleaved image-text scenarios with multiple images.
Based on this module, we propose MM-Interleaved, a novel end-to-end generation model for processing interleaved image-text data, as shown in Fig. 3. Images and text are first mapped into tokens through their respective tokenizers (i.e., image encoder and word embedding) and then fed to the LLM, arranged in their original order. A special token <BoI> is introduced to represent “Begin of Image”. When the LLM processes image and text sequences, each token in the intermediate layers directly observes multi-image and multi-scale feature maps from the previous context through the proposed feature synchronizer module. After being processed by the LLM, the features of the text tokens predict the next text word. When the <BoI> token is predicted, an image decoder based on diffusion models is used to predict the next image. All previous LLM output features are passed into the image decoder to generate images. Each pixel in the intermediate layers of the decoder can also extract details of the previous image through the proposed synchronizer module.
Our models are pre-trained on a mixture of image-text pairs and interleaved image-text sequences. We do not use any in-house data. Similar to previous multi-modal LLMs, supervised fine-tuning can further enhance our model capabilities. Thanks to end-to-end generative modeling, fine-tuning can be applied to both text and image generation. Our model is fine-tuned on several tasks, including visual question-answering, image caption, referring expression grounding, text-to-image generation, segmentation-to-image translation, and visual storytelling. Compared with previous image-to-text and text-to-image methods, our model achieves competitive results and shows higher token efficiency. Compared with joint image and text generation models, we set the new SotA results for a wide range of tasks. Overall, MM-Interleaved can handle interleaved image-text inputs and recognize fine-grained details across multiple images to produce accurate textual descriptions and visually consistent images, demonstrating the potential of high-performance multi-modal models that can understand and generate interleaved image-text content.
@@
Figure 2: Comparison of different types of image-text generative modeling. (a) and (b) can only generate text. (c) and (d) can generate both images and text. All types except (d) are limited by the fixed number of visual tokens extracted by context-insensitive Resampler, which can not efficiently capture image details and is problematic in multi-image scenarios.
@@ 2 Related Work
Modeling of Paired Image-Text Data plays an important role in the progress of multi-modal research in recent years. [37, 24, 76] has released a series of public large-scale image-text pairs datasets. Based on these data, models trained by image-text contrastive learning [70, 39, 82, 50] have been proven able to recognize and understand open-world semantics. Subsequent work [93, 111, 46, 97] further incorporates the text generation tasks such as image captioning. Some other works [74, 71] were proposed to generate images based on text prompts. The latest progress of LLMs [64, 65] has created a new era, leading to the emergence of many LLM-based multi-modal models trained using image-text pairs [48, 53, 90, 91]. For example, [53] connect pre-trained LLMs and vision foundation models with linear projections to build powerful multi-modal text-generation models.
Figure 3: Architecture of the proposed MM-Interleaved Model. The red lines represent how the multi-scale image features are generated and utilized. MM-Interleaved incorporates image encoder to both extract high-resolution multi-scale image features and map each image into a fixed number of low-resolution visual tokens. These visual tokens are fed into the multi-modal LLM along with text tokens. LLM uses the feature synchronization module to extract high resolution image details, and auto-regressively generates text tokens. After that, a DM-based image decoder generates next image conditioned on the previous context features from LLM as well as the multi-scale image features from the feature synchronization module.
@@
Modeling of Interleaved Image-Text Data has received increasing attention recently. [33, 3] were the first to focus on understanding such data relying on non-public datasets. To promote the development of this field, [109, 45] release public and large-scale interleaved image-text datasets. More works [103, 101] have begun to focus on understanding the interleaved data. Nevertheless, their generative capabilities remain limited to text. [43, 99, 83, 18] have initially attempted to generate images and text in the interleaved context concurrently. [43, 83] involve a two-stage generation process, where text and image are only generated in the first and second stages, respectively. [99] uses VQ-VAE [86] to convert images into discrete tokens, facilitating token-level auto-regressive modeling as traditional language modeling. However, it primarily focuses on image generation capabilities and exhibits notable weaknesses in image understanding. DreamLLM [18], a concurrent work to ours, also focuses on single-stage end-to-end modeling using raw image pixels as input. Despite these efforts, they only feed image information at the input of LLMs, which are limited by the problem that fixed number of visual tokens cannot efficiently describe image details.
@@
Integrating Image Details into LLMs is important for multi-modal LLMs. Most works use Perceiver Resamplers [3, 48, 107] to extract image information via cross-attention, mapping each image into a fixed number of visual tokens. For example, [3] equips Resamplers in the intermediate layers of LLMs, injecting extracted image features into LLMs through gated residual connections. [48, 107] introduces Resamplers at the bottom of LLM to insert the extracted visual tokens into the input text sequences. While these methods achieve good performance, image details may be overlooked due to the small number (e.g., 32 or 64) of visual tokens. To preserve more image details, [6, 55, 53, 10] increase the number of input visual tokens per image to hundreds. [84] further increases this number to 2,890. Although it mitigates information loss, the computational and memory demands of LLMs are significantly increased. Increasing the number of visual tokens per image is particularly problematic in multi-image scenarios, where multiple images naturally require more visual tokens, making it hard to use for interleaved image-text data.
@@ 3.1 Task Formulation
To build an end-to-end generative model for interleaved image-text data. We first consider an interleaved image-text sequence $X=\{x_{1},x_{2},x_{3},\dots\}$, where each element $x_{n}$ is either a text token (denoted as $x_{n}^{L}$) or a whole image (denoted as $x_{n}^{V}$). Text and images are arranged in the order in which they appear in the original content. In multi-modal LLMs, a common practice is first extracting embedding for each text token and each image before being fed into LLMs, i.e., $e_{n}^{L}=\mathcal{E}_{L}(x_{n}^{L})$ and $e_{n}^{V}=\mathcal{E}_{V}(x_{n}^{V})$, where $\mathcal{E}_{L}$ denotes word embedding following standard practices in NLP. $\mathcal{E}_{V}$ is typically an image encoder (e.g., ViTs [19]) followed by a Perceiver Resampler [3] to map each image to a fixed number of tokens. Then, the generative modeling is trained to maximize the log-likelihood of the multi-modal sequence, given by:
$\log p(X)$ $=\sum_{n}\log p(x_{n}|x_{1},x_{2},\dots,x_{n-1})$ (1)
$=\sum_{n}\log p(x_{n}|e_{1},e_{2},\dots,e_{n-1})$
$=\sum_{n\in\mathcal{I}_{L}}\log p(x_{n}^{L}|e_{<n})+\sum_{n\in\mathcal{I}_{V}}\log p(x_{n}^{V}|e_{<n}),$
@@
where $\mathcal{I}_{L}$ and $\mathcal{I}_{V}$ represent the index sets for text tokens and images within the sequence, respectively. $<n$ in the subscript represents the abbreviation of $\{1,2,\dots,n-1\}$. The following paragraphs provide explanations of Eq. (1).
@@
Text Generation with Multi-modal Condition. The log-likelihood $\log p(x_{n}^{L}|e_{<n})$ is similar to traditional causal language modeling, except that the condition also includes previous images. Recent works [48, 53, 90, 91] have demonstrated the effectiveness of using LLMs for processing additional visual inputs. They apply a linear classifier on top of the LLMs. The loss function for text generation is
$L_{\text{NTP}}(x_{n}^{L}|e_{<n})=-\bar{x}_{n}^{L}\cdot\log\mathrm{softmax}\big{(}W\cdot\mathcal{D}_{\text{LM}}(e_{<n})\big{)},$ (2)
@@
where $W$ is the linear classification weight, $\mathcal{D}_{\text{LM}}$ denotes the LLM network (e.g., LLaMA [85]), $\bar{x}_{n}^{L}$ is the one-hot vector indicating the ground-truth text token.
@@
Image Generation with Multi-modal Condition. Maximizing the log-likelihood $\log p(x_{n}^{V}|e_{<n})$ aligns with the denoising-diffusion process [54], which recently achieves widespread success in image generation. When conditional inputs $x_{<n}$ are all text, this is identical to text-to-image diffusion models, where maximizing the log-likelihood is derived as minimizing the diffusion modeling loss as
$L_{\text{NIP}}(x_{n}^{V}|e_{<n})=\mathbb{E}_{\epsilon,t}~{}||\epsilon-\mathcal{D}_{\text{DM}}\big{(}x_{n,t}^{V},t,\mathcal{D}_{\text{LM}}(e_{<n})\big{)}||^{2},$ (3)
@@
where $\mathcal{D}_{LM}$ is the language model encoding text inputs, $\mathcal{D}_{DM}$ is the diffusion model for image denoisiong. $x_{n,t}^{V}$ is the noisy version of the original image at the denoising step $t$, and the denoising network $\mathcal{D}_{DM}$ is trained to predict the noise $\epsilon$. We found that such modeling is also applicable when conditional input is interleaved image-text data.
@@
Note that our modeling allows for the flexible combination of different language models, image encoders, and image decoders to fully leverage a variety of pre-trained models. It does not need to insert additional special tokens specifically for image generation into LLMs. The whole framework can be optimized end-to-end.

### v2

@@ Abstract
* Equal contribution; 🖂 Corresponding author (daijifeng@tsinghua.edu.cn).† This work is done when Changyao Tian, Weiyun Wang, and Zhe Chen are interns at Shanghai AI Laboratory.
Developing generative models for interleaved image-text data holds both research and practical value. It necessitates models to comprehend interleaved sequences and subsequently generate images and text. However, existing attempts are limited by the issue that the fixed number of visual tokens cannot efficiently capture image details, which is particularly problematic in the multi-image scenarios. To address this, this paper presents MM-Interleaved, an end-to-end generative model for interleaved image-text data. It introduces a multi-scale and multi-image feature synchronizer module (MMFS), enabling intermediately direct access to fine-grained image features from the previous context during the generation process. MM-Interleaved is end-to-end pre-trained on both paired and interleaved image-text corpora. It is further enhanced through a supervised fine-tuning phase, wherein the model improves its ability to follow complex multi-modal instructions. Experiments demonstrate the versatility of MM-Interleaved in recognizing visual details following multi-modal instructions and generating consistent images following both textual and visual conditions. Code and models are available at https://github.com/OpenGVLab/MM-Interleaved.
Keywords:
Interleaved image-text modeling Large language models
@@ 1 Introduction
Interleaved image-text data like news and blogs is a sequence of multiple images interspersed with text, which is ubiquitous on the internet. As an extension of image-text pairs widely used in previous multi-modal models [37, 21, 83, 20, 71], the interleaved format not only covers a broader range of data but also presents longer and more complex article structures. Developing multi-modal models capable of simultaneously comprehending and generating such interleaved image-text data has significant research value and practical application potential. It expands the scope of previous multi-modal models [84] and enables more unified processing of multi-modal data, bridging previously disparate research fields such as text-to-image generation [74, 73] and visual question-answering [47, 54].
@@
With recent progress in multi-modal modeling with Large Language Models (LLMs) [7, 65], exploring how LLMs can be leveraged for interleaved image-text modeling has become increasingly attractive. A core challenge of interleaved multi-modal LLMs is how to effectively handle multiple images within the limited context length of LLMs. In most multi-modal LLMs, images are encoded as visual tokens from an image tokenizer (also known as image encoder) and fed into LLMs together with text tokens [43, 17, 84, 40]. However, due to computational and memory constraints, multi-modal LLMs have limited context window sizes (e.g., 2048 or 4096 tokens). To reduce computational demands and the required context length of LLMs, Perceiver Resamplers [3] are commonly used to map each image from up to 1024 tokens to a small fixed number of visual tokens (e.g., 32 or 64 tokens) as in Fig. 2.
Due to the relatively small visual token number, critical image details may be overlooked, especially in tasks requiring fine-grained observation. Although increasing the number of visual tokens [56, 101] may address this issue to some extent, the token number per image must be restricted to ensure complete input into the LLMs. Allocating a large number (e.g., 2890 in [85]) of visual tokens per image would pose a significant limitation especially in multi-image scenarios.
Figure 1: Illustration of Multi-Modal Feature Synchronizer (MMFS). In the auto-regressive generation of interleaved image-text sequences, besides interacting with low-resolution image tokens via self-attention, tokens in MM-Interleaved could also use MMFS to cross-attend multi-scale high-resolution image features. MMFS ensures the attention will have exactly the same causal relation between images and text tokens.
@@
We note that the main problem here is the context insensitivity of Perceiver Resamplers, as it only takes image features as input. Each input image is compressed into a fixed number of tokens, making it challenging or even infeasible to preserve all requisite information and accommodate subsequent generation requirements. However, such information loss can actually be mitigated through further feature extraction in the intermediate layers of the LLMs. Based on the intermediate features of the LLMs, relevant information can be dynamically extracted from the original images to fulfill the perceptual requirements. Such a dynamic extraction mechanism can also help to effectively handle scenarios involving multiple images and high-resolution image feature maps.
@@
Figure 2: Different types of image-text generative modeling. (a) and (b) can only generate text. (c) and (d) can generate both images and text. All types except (d) are limited by the fixed number of visual tokens extracted by context-insensitive Resampler, which will lose image details and is problematic in multi-image scenarios.
To facilitate dynamic image feature extraction in the intermediate layers of the LLMs, we design a fine-grained Multi-Modal Feature Synchronizer (MMFS) based on the Deformable Sparse Attention [112], which can efficiently extract relevant information from multi-scale image feature maps and multiple images as shown in Fig. 1. Based on MMFS, we further propose MM-Interleaved, a novel end-to-end generation model for processing interleaved image-text data, as shown in Fig. 4, which can generate both textual and visual response given interleaved image-text inputs.
Specifically, input images and text are first mapped into tokens through their respective tokenizers (i.e., image encoder and word embedding) and then fed to the LLM, arranged in their original order. A special token <BoI> is introduced to represent “Begin of Image”. When the LLM processes the input sequences, each token in the intermediate layers directly observes multi-image and multi-scale feature maps from the previous context through MMFS. After being processed by the LLM, the features of the text tokens are then used to predict the next text word. When the <BoI> token is predicted, an image decoder based on diffusion models is used to predict the next image. All previous LLM output features are passed into the image decoder as the generation conditions. With MMFS, the image decoder can also extract details of previous images intermediately.
MM-Interleaved is pre-trained on a mixture of image-text pairs and interleaved image-text sequences without using any in-house data. Similar to previous multi-modal LLMs, supervised fine-tuning can further enhance the model capabilities. Due to the end-to-end generative modeling, fine-tuning can be applied to both text and image generation. Our model is fine-tuned on several tasks, including visual question-answering, image captioning, referring expression grounding, text-to-image generation, segmentation-to-image translation, and visual storytelling. As is illustrated in Fig. 3, our model achieves competitive results and shows higher token efficiency compared with previous image-to-text and text-to-image methods. Compared with joint image and text generation models, we set the new SOTA results for a wide range of tasks. The key contributions are:
  • We propose Multi-Modal Feature Synchronizer (MMFS) to reduce the number of visual tokens required by multi-modal LLMs, which can efficiently extract fine-grained visual details from multi-scale feature maps of multiple images according to the intermediate context features of the multi-modal LLMs;
  • We propose MM-Interleaved based on MMFS for generative modeling of interleaved image-text data, which can preserve fine-grained image information with only a small number of visual tokens per image, and can be optimized end-to-end for both text and image generation;
  • Our method can generate both accurate text descriptions and visually consistent images given interleaved image-text inputs, achieving SOTA results on various multi-modal comprehension and generation benchmarks without using any in-house data.
Figure 3: Comparison on multi-modal benchmarks. 11 representative tasks are listed. See Sec. 4 for other benchmarks. MM-Interleaved (red) consistently outperforms previous methods supporting interleaved generation (green). It also surpasses or is on par with specialists in either text generation (brown) or image generation (blue).
@@ 2 Related Work
Modeling of Paired Image-Text Data plays an important role in the progress of multi-modal research in recent years. A series of public large-scale image-text pairs datasets have been released [37, 23, 77]. Based on these data, models trained by image-text contrastive learning [71, 39, 83, 50] are able to recognize and understand open-world semantics. Subsequent works [94, 113, 49, 99] further incorporated text generation tasks such as image captioning, while other works [74, 73] were proposed to generate images based on text prompts. The latest progress of LLMs [65, 66] has launched a new era, leading to the emergence of many LLM-based multi-modal models trained using image-text pairs [47, 54, 92, 91]. For example, LLaVA [54] connects LLMs with pretrained visual encoders using linear projections to build powerful multi-modal foundation models.
@@
Modeling of Interleaved Image-Text Data has received increasing attention recently. Early works such as Kosmos [33] and Flamingo [3] focused on understanding such data with non-public datasets. To promote the development of this field, public and large-scale interleaved image-text datasets were later released in [111, 45]. More recent works [105, 103] have concentrated on understanding the interleaved data. Nevertheless, their generative capabilities remain confined to text. Initial attempts at generating images and text given interleaved contexts were first undertaken in [43, 101, 84, 17]. A two-stage generation process was introduced by [43, 84], wherein text and images are generated in the first and second stages, respectively. CM3Leon [101] employs VQ-VAE [87] to convert images into discrete tokens, facilitating token-level auto-regressive modeling as language modeling. However, it primarily focuses on image generation capabilities and exhibits notable weaknesses in image understanding. DreamLLM [17] focuses on single-stage end-to-end modeling using raw image pixels as inputs. Emu2 [82], SEED-LLaMA [25] and VL-GPT [110] adopt additional training stage for the image tokenizer-detokenizer. Despite these efforts, they only feed image information at the LLM inputs, which are limited by the problem that fixed number of visual tokens cannot efficiently preserve image details.
@@
Integrating Image Details into LLMs is important for multi-modal LLMs. Most works use Perceiver Resamplers [3, 47, 109] to extract image information via cross-attention, mapping each image into a fixed number of visual tokens. For example, Flamingo [3] adopts Resamplers in the intermediate layers of LLMs, injecting extracted image features into LLMs through gated residual connections. BLIP-2[47] and Mini-GPT4 [109] introduce Resamplers at the bottom of LLM to insert the extracted visual tokens into the input text sequences. While these methods achieve good performance, image details remain to be overlooked due to the small number (e.g., 32 or 64) of visual tokens. To preserve more details, recent works [6, 56, 54, 10] increase the number of input visual tokens per image to hundreds. In SPHINX [85], the token number is further increased to 2,890. Although it helps mitigate information loss, the computational and memory demands of LLMs are also significantly increased. Increasing the number of visual tokens per image is particularly problematic in multi-image scenarios, where multiple images naturally require more visual tokens, making it hard to use for interleaved image-text data.
@@ 3.1 Task Formulation
To build an end-to-end generative model for interleaved image-text data, we first consider an interleaved image-text sequence $X=\{x_{1},x_{2},x_{3},\dots\}$, where each element $x_{n}$ is either a text token (denoted as $x_{n}^{L}$) or a whole image (denoted as $x_{n}^{V}$). Text and images are arranged in the order in which they appear in the original content. In multi-modal LLMs, a common practice is to first extract embeddings for each text token and each image and then feed them into LLMs, i.e., $e_{n}^{L}=\mathcal{E}_{L}(x_{n}^{L})$ and $e_{n}^{V}=\mathcal{E}_{V}(x_{n}^{V})$, where $\mathcal{E}_{L}$ denotes word embedding following standard practices in NLP. $\mathcal{E}_{V}$ is typically an image encoder (e.g., ViTs [18]) followed by a Perceiver Resampler [3] to map each image to a fixed number of tokens. Then, the generative modeling is trained to maximize the log-likelihood:
$\log p(X)=\sum_{n}\log p(x_{n}|e_{<n})=\sum_{n\in\mathcal{I}_{L}}\log p(x_{n}^{L}|e_{<n})+\sum_{n\in\mathcal{I}_{V}}\log p(x_{n}^{V}|e_{<n}),$ (1)
@@
where $\mathcal{I}_{L}$ and $\mathcal{I}_{V}$ represent the index sets for text tokens and images, respectively. $<n$ in the subscript represents the abbreviation of $\{1,2,\dots,n-1\}$. The following paragraphs provide explanations of Eq. (1).
Figure 4: Architecture of MM-Interleaved. The red lines represent how the multi-scale image features are generated and utilized. MM-Interleaved incorporates image encoder to both extract high-resolution multi-scale image features (red lines) and map each image into a fixed number of low resolution visual tokens. These visual tokens are fed into the multi-modal LLM along with text tokens. LLM then uses our proposed feature synchronizer (MMFS) to extract additional high-resolution image details, and auto-regressively generate text tokens. After that, a diffusion-based image decoder generates next image conditioned on the previous context features from LLM, where MMFS is also utilized to capture accurate visual conditions.
@@
Text Generation with Multi-modal Condition. $\log p(x_{n}^{L}|e_{<n})$ is similar to traditional causal language modeling, except that the condition also includes previous images. Recent works [47, 54, 92, 91] have demonstrated the effectiveness of using LLMs for processing additional visual inputs. They apply a linear classifier on top of the LLMs. The loss function for text generation is
$L_{\text{NTP}}(x_{n}^{L}|e_{<n})=-\bar{x}_{n}^{L}\cdot\log\mathrm{softmax}\big{(}W\cdot\mathcal{D}_{\text{LLM}}(e_{<n})\big{)},$ (2)
@@
where $W$ is the linear classification weight, $\mathcal{D}_{\text{LLM}}$ denotes the LLM network (e.g., LLaMA [86]), $\bar{x}_{n}^{L}$ is the one-hot vector indicating the ground-truth text.
@@
Image Generation with Multi-modal Condition. Maximizing $\log p(x_{n}^{V}|e_{<n})$ aligns with the denoising-diffusion process [55], which recently achieves widespread success in image generation. Maximizing the log-likelihood is derived as minimizing the diffusion modeling loss as
$L_{\text{NIP}}(x_{n}^{V}|e_{<n})=\mathbb{E}_{\epsilon,t}~{}||\epsilon-\mathcal{D}_{\text{DM}}\big{(}x_{n,t}^{V},t,\mathcal{D}_{\text{LLM}}(e_{<n})\big{)}||^{2},$ (3)
@@
where $\mathcal{D}_{DM}$ is the diffusion model for image denoisiong. $x_{n,t}^{V}$ is the noisy version of the original image at the denoising step $t$, and the denoising network $\mathcal{D}_{DM}$ is trained to predict the noise $\epsilon$. We found that such modeling is also applicable when conditional input is interleaved image-text data.
@@
Such modeling allows for the flexible combination of different language models, image encoders, and image decoders to fully leverage a variety of pre-trained models. The entire framework can be optimized end-to-end.